DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, May 27, 2004 Padelford C-36 1:30 pm Geometric view points for estimates of elliptic equations

Lihe Wang (University of Iowa)

We will talk about L^{p} estimates for elliptic and parabollic equations for domains that are geometrically invariant.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.